If the perimeter of the triange is 32 inches, what is the length of the longest side if there are two equal sides and a third side that is five inches more than the length of the two equal sides?
Answer by rfer(16322)

x+x+x+5=32
3x=27
x=27/9
x=9
x+5=14 long side